OpenStudy (anonymous):

Let f(x) = 3x+2 and g(x) = x2 Part 1 [2 points] Find f(g(x)) Part 2 [2 points] Find g(f(x)) Part 3 [4 points] Use complete sentences to explain any difference in part 1 and part 2. PLEASE HELP IM STUMPED

OpenStudy (anonymous):

compose the functions within each other f(g(x)) = f(x^2) = 3x^2+2 g(f(x)) = g(3x+2) = (3x+2)^2

OpenStudy (anonymous):

part 1 \[f(g(x))=3*g(x)+2=3(x2)+2=3x2+2\]